If 3a + 4b=b which of the following must equal 6a + 6b?
(a). 0
(b). 12
(c) 12b

Answer by nerdybill(7384)

3a + 4b=b
subtracting b from both sides:
3a + 3b=0
multiplying both sides by 2:
6a + 6b=0
.
Answer: 'a'